# QG-SMS: Enhancing Test Item Analysis via Student Modeling and Simulation

### v1

@@ 6 Related Work
Student Modeling and Simulation with (L)LMs: Recent studies explore the use of (L)LMs to simulate human behaviors in general Park et al. (2023), and classroom learning in particular Xu and Zhang (2023); Zhang et al. (2024). These simulations have been applied in various educational contexts, from training novice teachers Markel et al. (2023) to promoting student engagement Zhang et al. (2024). Prior works have utilized LM-based simulations for evaluating test items, with Park et al. (2024) and Moon et al. (2022) using multiple (L)LMs with varying capacities to model different students in the classroom for assessing question answerability and difficulty. However, these studies overlook key dimensions such as item discrimination and distractor efficiency. Unlike these approaches, our proposed method, QG-SMS, demonstrates that a single LLM is capable of simulating students at diverse levels, making the pipeline more efficient and scalable. While Lu and Wang (2024) manually specify knowledge mastery levels in the prompt to the LLM, our approach eliminates this need, making simulation more flexible. Additionally, we conduct comprehensive experiments to further validate the usefulness of simulated student profiles for test item analysis.
@@ 7 Conclusion
In this work, we proposed QG-SMS, a novel simulation-based QG evaluation framework for test item analysis. We first constructed two datasets of candidate question pairs that differ in quality across multiple dimensions of educational value. Experiments with existing evaluation approaches highlight the challenges of accurately and efficiently assessing test item quality. In response, we introduce the modeling and simulation of diverse student understanding for evaluation. These simulated student profiles offer valuable insights into how well a question functions as a test item for assessing student performance. We conducted experiments across two datasets and four dimensions of test item analysis, as well as recruited human annotators to showcase the effectiveness, robustness, and adaptability of QG-SMS in performing comprehensive test item analysis.
@@
In this work, we evaluate the quality of test items at an individual level. We recognize that constructing assessment typically requires considering multiple dimensions and ensuring diversity within each dimension Osterlind (1997). For example, a well-designed quiz should not only cover different topics from the learning materials rather than repeatedly assessing the same concept, but also include a mix of easy, medium, and hard questions. One potential application of QG-SMS in such scenarios is to rank candidate test items based on a given dimension $d$ by comparing simulated student understanding and performance. Using these rankings, future work could explore methods to assist teachers in assembling assessments that achieve balance across relevant dimensions.
@@ A.1 Prompts for QG-SMS
Figure 6: Prompts for our three-step evaluation approach QG-SMS.
@@ A.2 Experimental Details
Underlying LLM: For all LLM-based experiments with GPT-4o, we used the gpt-4o-2024-05-13 checkpoint.
Baseline implementation: For BERTScore, we use the implementation of Hugging Face evaluate [^2] https://huggingface.co/docs/evaluate/en/index package (bertscore). For KDA [^3] https://github.com/riiid/question-score and ChatEval  [^4] https://github.com/thunlp/ChatEval, we used the code implementation provided by the authors. To obtain the expert personas for ChatEval, we utilized the AutoAgents interactive framework [^5] https://github.com/Link-AGI/AutoAgents given instruction $I$ as described in §$\S$4.2. We used the implementation by Zeng et al. 2024666https://github.com/princeton-nlp/LLMBar for the remaining LLM-based evaluation approaches.
@@ A.3 Human Evaluation Details
Selection of human-written question pairs: In the EduAgent dataset, both questions in a ($Q_{1}$, $Q_{2}$) pair comes from the same lecture. However, they can be grounded to either the same or different sections of the lecture. For example, in Fig. 1, $Q_{1}$ is relevant to the Introduction to computer vision section, while $Q_{2}$ is relevant to the Computer vision history section. To reduce the cognitive load for annotators, we opt for question pairs that are grounded to the same section in the same lecture. Based on this condition, we selected 60 pairs of human-written questions that exhibit differing quality: 21 pairs in the DF dimension, 15 pairs in the DC dimension, and 24 pairs in the DE dimension.

### v2

@@ 6 Related Work
Student Modeling and Simulation with (L)LMs: Recent studies explore the use of (L)LMs to simulate human behaviors in general Park et al. (2023), and classroom learning in particular Xu and Zhang (2023); Zhang et al. (2024). These simulations have been applied in various educational contexts, from training novice teachers Markel et al. (2023) to promoting student engagement Zhang et al. (2024). Prior works have utilized LM-based simulations for evaluating test items. Some limit the simulation to a single group of students Säuberli and Clematide (2024), while others use multiple (L)LMs with varying capacities to model different students in the classroom Lalor et al. (2019); Moon et al. (2022); Park et al. (2024). Unlike these approaches, our proposed method demonstrates that a single LLM is capable of simulating students at diverse levels, making the pipeline more efficient and scalable. While the approaches proposed by Lu and Wang (2024); Lalor et al. (2019); Hayakawa and Saggion (2024); Byrd and Srivastava (2022) require manual efforts to control the simulated student profiles through either feature engineering or prompt engineering, our approach eliminates this need, making simulation more flexible.
@@ 7 Conclusion
In this work, we proposed QG-SMS, a novel simulation-based QG evaluation framework for test item analysis. We first constructed two datasets of candidate question pairs that differ in quality across multiple dimensions of educational value. Experiments with existing evaluation approaches highlight the challenges of accurately and efficiently assessing test item quality. In response, we introduced the modeling and simulation of diverse student understanding for evaluation. These simulated student profiles offer valuable insights into how well a question functions as a test item for assessing student performance.
We identify two promising future directions. Prior work has shown that, despite being prompted with educational requirements, LLMs often fail to incorporate them into generated questions, as judged by human evaluators Al Faraby et al. (2024). We have shown that QG-SMS is a reliable indicator of educational aspects like DF, DE, and DC. In this sense, QG-SMS could be integrated into a reward-based optimization pipeline to better align generated test items with educational objectives. Additionally, we observe a growing interest in research question generation Liu et al. (2024a, b), which will potentially benefit from a simulation-based evaluation framework like QG-SMS. Existing works still rely on costly and time-consuming evaluation by human researchers. Future work could explore simulating diverse researcher perspectives to enable automated, scalable evaluation of research questions.
@@
In this work, we evaluate the quality of test items at an individual level. We recognize that constructing assessment typically requires considering multiple dimensions and ensuring diversity within each dimension Osterlind (1997). For example, a well-designed quiz should not only cover different topics from the learning materials rather than repeatedly assessing the same concept, but also include a mix of easy, medium, and hard questions. One potential application of QG-SMS in such scenarios is to rank candidate test items based on a given dimension $d$ by comparing simulated student understanding and performance. Using these rankings, future work could explore methods to assist teachers in assembling assessments that achieve balance across relevant dimensions. Additionally, our significance tests rely on a p-value threshold of $0.1$ (see Appendix A.2 for more details). Future works could explore whether stronger models could lead to more robust significance results.
@@ A.1 Prompts for QG-SMS
Step 1: Student Profile Generation Given the following learning materials: {Lecture Content / Knowledge Component Descriptions $L$} Consider students with various understanding in a scenario where a quiz about the above learning materials is being conducted. Ensure that you generate at least 10 roles for the scenario. For each student, provide a detailed description that includes their name and their understanding of the lecture content. The distribution of understanding of lecture content must mimic that in a real classroom.
Step 2: Student Performance Prediction Given the following learning materials: {Lecture Content / Knowledge Component Descriptions $L$} Below is the list of students and their reported understanding of the learning materials: {Student Profiles from Step 1} Given the following quiz questions about the lecture content: Question 1: {Question $Q_{1}$} Question 2: {Question $Q_{2}$} For each student, predict whether the student will correctly answer each question based on both the student’s understanding, question’s difficulty, guessing factors, etc.). If you predict “incorrect”, specify which distractor confuses the student.
Step 3: Evaluation You are interested in finding a quiz question that satisfies the following requirement: {Requirement $R_{d}$} You are given 2 output quiz questions Output (a) and Output (b) and the analysis of the responses of each student who attempted the questions. Your task is to identify which of Output (a) and Output (b) better satisfies requirement {$R_{d}$} based on the question content and student performance. {Description of $R_{d}$} # Output (a): {Question $Q_{1}$} # Output (b): {Question $Q_{2}$} # Consider Students Performance: {Predicted student performance from Step 2} # Which question better satisfies {$R_{d}$}, Output (a) or Output (b)? Your response should be either "Output (a)" or "Output (b)"
Figure 5: Prompts for our three-step evaluation approach QG-SMS.
@@ A.2 Experimental Details
Construction of TC pairs: For both datasets, each question is annotated with one or more related topics. From all possible pairs in the dataset, we first perform filtering. For EduAgent, each question is associated with a specific section of a lecture. We only consider pairs of questions from the same section of the same video lecture. The learning material $L$ consists of the lecture transcript and slide descriptions. For DBE-KT, each question may be linked to one or more knowledge components. We only consider pairs with exactly one differing knowledge component and at least one shared component, making the test cases more challenging. Here, the learning material L is the union of the knowledge components for the pair. Then, for each selected pair, we randomly choose one question as the label (TC = 1). The associated topic is set as the knowledge components (DBE-KT) or section (EduAgent) of this label. The other question in the pair is considered dispreferred output (TC = 0). This set up results in 286 pairs for DBE-KT and 217 pairs for EduAgent.
Number of generated student profiles: We are inspired by Pang et al. (2024), which simulates social scenarios for LLM alignment and selects 10 generated profiles to balance diversity (crucial for evaluating question quality) and computational efficiency. Following this approach, our prompt instructs the model to generate at least 10 roles for each learning material L. We then use all generated profiles—without filtering or augmentation—for Steps 2 and 3.
Underlying LLM: For all LLM-based experiments with GPT-4o, we used the gpt-4o-2024-05-13 checkpoint with the default hyperparameters. Regarding the experiments on the robustness of the student profiles (§5.2), we ran the same prompt for Step 1 multiple times using the same default hyperparameters (temperature $=1$). Thus, any differences in output come from the random seeds used in the API calls.
Baseline implementation: For BERTScore, we use the implementation of Hugging Face evaluate [^3] https://huggingface.co/docs/evaluate/en/index package (bertscore). For KDA [^4] https://github.com/riiid/question-score, ChatEval  [^5] https://github.com/thunlp/ChatEval, and QSalience [^6] https://github.com/ritikamangla/QSalience/, we used the code implementation provided by the authors. To obtain the expert personas for ChatEval, we utilized the AutoAgents interactive framework [^7] https://github.com/Link-AGI/AutoAgents given instruction $I$ as described in §$\S$4.2. We used the implementation by Zeng et al. 2024888https://github.com/princeton-nlp/LLMBar for the remaining LLM-based evaluation approaches.
Comparison direction for single-scoring metrics
  • Higher Topic Alignment: $\uparrow$ BERTScore, $\uparrow$ KDA, $\downarrow$ QSalience
  • Easier question: $\downarrow$ BERTScore, $\uparrow$ KDA, $\downarrow$ QSalience
  • Higher discrimination: $\uparrow$ BERTScore, $\downarrow$ KDA, $\uparrow$ QSalience
  • Higher distractor efficiency:$\uparrow$ BERTScore, $\downarrow$ KDA, $\uparrow$ QSalience
Signficance testing: We consider Vanilla, which simply outputs a preference given the instruction without any reasoning or intermediate steps, the base strategy for using LLMs in test item analysis. Other baselines and QG-SMS can be considered more complex strategies for using LLMs as evaluators of test items. We conduct pairwise binomial tests to examine whether each LLM-based evaluation approach (including QG-SMS) significantly improves consistent accuracy in test item analysis compared to Vanilla. We report the p-values of the binomial tests on QG-SMS improvements over the base strategy (Vanilla), as compared to all other evaluation approaches in Tbl. 6.
Method DF DC DE
EduAgent DBE-KT EduAgent DBE-KT EduAgent
CoT 0.999 0.996 0.999 0.999 0.999
Metrics 0.227 0.640 0.500 0.910 0.813
Reference 0.500 0.925 0.938 0.967 0.981
Swap 0.192 0.134 0.856 0.588 0.985
ChatEval 0.500 0.007 0.927 0.262 0.942
QG-SMS 0.007 0.000 0.252 0.124 0.093
Table 6: P-values from binomial tests assessing whether the LLM-based evaluation strategy significantly improves consistent accuracy compared to the Vanilla baseline.
@@ A.3 Human Evaluation Details
Selection of human-written question pairs: In the EduAgent dataset, both questions in a ($Q_{1}$, $Q_{2}$) pair comes from the same lecture. However, they can be grounded to either the same or different sections of the lecture. For example, in Tbl. 1, $Q_{1}$ is relevant to the Introduction to computer vision section, while $Q_{2}$ is relevant to the Computer vision history section. To reduce the cognitive load for annotators, we opt for question pairs that are grounded to the same section in the same lecture. Based on this condition, we selected 60 pairs of human-written questions that exhibit differing quality: 21 pairs in the DF dimension, 15 pairs in the DC dimension, and 24 pairs in the DE dimension.
@@
Difficulty-controlled question generation Given the following learning materials: {Lecture Content / Knowledge Component Descriptions $L$} Generate 4-choice quiz questions to test students’ understanding of the lecture. The generated questions should have diverse difficulty. The more difficult a question, the fewer number of students can correctly answer it. There must be 2 (two) ’easy-level’ questions, 2 (two) ’medium-level’ questions, and 2 (two) ’hard-level’ questions.
Discrimination-controlled question generation Given the following learning materials: {Lecture Content / Knowledge Component Descriptions $L$} Generate 4-choice quiz questions to test students’ understanding of the learning materials. The generated questions should have diverse discrimination ability. A question with high discrimination is more effective at distinguishing between high-performing and low-performing students. An example of a question with low discrimination is when neither high-performing nor low-performing students can answer the question correctly, or when all students can answer the question correctly. There must be 2 (two) ’low-discrimination’ questions, 2 (two) medium-discrimination questions, and 2 (two) ’high-discrimination’ questions.
Distractor-efficiency-controlled question generation Given the following learning materials: {Lecture Content / Knowledge Component Descriptions $L$} Generate 4-choice quiz questions to test students’ understanding of the lecture. The generated questions should have diverse number of effective distractors. An effective distractor is one that will be selected by at least 5% of the students. Specifically, there must be 2 (two) questions with NO effective distractors, 2 (two) questions with exactly ONE effective distractors, 2 (two) questions with exactly TWO effective distractors, and 2 (two) questions with all THREE effective distractors.
@@
A.4 Qualitative Analysis on Limitations of Existing Evaluation Approaches
We provide in Tbl. 8 the generations of all LLM-based baselines when evaluating the two questions from the case study presented in Tbl. 1. As shown, strategies that incorporate reasoning (e.g., ChatEval, CoT, Swap) consistently prioritize questions that “apply” the concept of computer vision—indicating higher discrimination—over questions that require recalling specific statistics. Meanwhile, generation-based criteria (Metrics, Reference) are too broad and not specific enough to assess the requirement of high discrimination, likely due to being distracted by the long learning material context. Without explicitly modeling student performance, it is difficult for LLMs to "reason" about discrimination based solely on question content and materials.
Learning Material
Introduction of computer vision: Computer vision (CV) is the field of computer science that focuses on creating digital systems that can process, analyze, and make sense of visual data […]. For example, […]
Computer vision history […] In 2012, a team from the University of Toronto […].
The model, called AlexNet, […], achieved an error rate of 16.4%, which overperformed all other methods at that time. […]
Quiz Questions
$Q_{1}$: Which of the following may utilize computer vision techniques? 1). Use a camera to check potential issues on the surface of products (2). Estimate the freshness of apples from pictures (3). Estimate whether a car is speeding via a camera (4). Determine whether a piece of audio is spoken by a specific person
A) (1)(2)(3); B) (1)(2)(4); C) (2)(3)(4); D) (1)(2)(3)(4).
$Q_{2}$: One breakthrough in computer vision happened at the University of Toronto in 2012, which achieved an error rate of [ ] in image classification.
A) 6.4%; B) 10.4% C) 12.4% D) 16.4%.
Evaluation Task: Which question has higher discrimination?
Existing approaches: $Q_{1}$. $Q_{1}$ is an apply-level question, while $Q_{2}$ is a recall-level question.
Label based on Actual Student Performance: $Q_{2}$. Applications of CV appearing in $Q_{1}$ can be considered common knowledge while $Q_{2}$ tests a specific detail which only students who pay close attention to details may be able to answer.
CoT: Q1 asks students to identify which scenarios utilize computer vision techniques, which requires a deeper understanding of the lecture content and the application of computer vision. This question effectively distinguishes between students who have a comprehensive understanding of the lecture and those who do not, as it requires them to apply their knowledge to different scenarios. Q2, on the other hand, asks for a specific fact from the lecture, the error rate achieved by the University of Toronto in 2012. While this question tests recall, it does not require the same level of understanding or application of knowledge as Q1. It is more about memorizing a specific detail rather than understanding the broader concepts and applications of computer vision.
Metrics: Does the question accurately and comprehensively reflect the specific content and structure of the lecture slides and transcripts provided in the instruction? Does the question maintain the clarity and coherence of the information as presented in the lecture slides and transcripts? Does the question include all the required elements mentioned in the instruction without adding or omitting any details?
Reference: Explain the significance of the 2012 breakthrough by the University of Toronto in the context of computer vision history, and discuss the two key factors that contributed to the success of convolutional neural networks (CNNs) at that time.
Swap: Q1 presents a question that requires students to apply their understanding of computer vision applications, which were discussed in the lecture. This question effectively tests the students’ ability to distinguish between tasks that can and cannot be performed using computer vision, thus having high discrimination. Q2 focuses on a specific fact mentioned in the lecture, the error rate achieved by the University of Toronto in 2012. While this question tests recall of a specific detail, it does not effectively distinguish between high-performing and low-performing students as it relies more on memorization rather than understanding and application of concepts.
ChatEval $Q_{1}$: Relevance to Lecture Content: The question is relevant as it pertains to the applications of computer vision, which is discussed in the lecture. Complexity and Depth of Knowledge: The question requires students to apply their understanding of computer vision to different scenarios, which is a higher-order cognitive skill. This makes it effective at distinguishing between high-performing and low-performing students. Appropriateness: The question is well-designed but includes an option related to audio, which is not directly related to computer vision. This could potentially confuse students.
$Q_{2}$: Relevance to Lecture Content: The question is directly relevant to the lecture content, specifically the historical breakthrough in computer vision. Complexity and Depth of Knowledge: The question tests recall of a specific fact, which is a lower-order cognitive skill. It does not require deep understanding or application of knowledge, making it less effective at distinguishing between high-performing and low-performing students. Appropriateness: The question is straightforward and tests a specific detail from the lecture, but it lacks complexity and does not challenge students to apply their knowledge.
Table 8: Case Study illustrates limitations of existing evaluation metrics in test item analysis.